OpenStudy (anonymous):

Okay, this question had me and my friend confused today. In a certain electrical circuit, the current "i" satisfies the differential equation:- L (di / dt) = E - Ri (a) Solve the equation to show that E = Ri + Ae^(-Ri / L)

OpenStudy (jamesj):

The solution you have written down isn't quite right. But in any case, this solution is easily solved by what is called separation of variables: \[ L \frac{di}{dt} = E - Ri \ \ \Longrightarrow \ L \frac{di}{E-Ri} = dt\] Now integrate... \[ \int L \frac{di}{E-Ri} = \int dt \] Hence ... you take it from here

OpenStudy (anonymous):

Oh sorry, the solution should be "E = Ri + Ae^(-Rt / L)" :) Would L in this case be taken over to the other side to get \[\int\limits \frac{ di }{ E - Ri } = \int\limits \frac{ dt }{ L }\] From here I ended up with \[\ln |E - Ri| = \frac{ t }{ L }\] Correct me if im wrong ;o

OpenStudy (anonymous):

\[\ln |E = Ri| = \frac{ t }{ L } + C\] I mean, can't miss out the C ;)

OpenStudy (jamesj):

Not quite. You have a -R in front of the i, so when you integrate, we had better see that appear somewhere as constant.

OpenStudy (anonymous):

Oh I see :) thanks.